(D

3 − 3DD

′2 − 2D

′3

)z = cos(x + 2y)


1
Expert's answer
2021-09-16T08:05:28-0400

(D3-3DD'-2D'3)z=Cos(x+2y)

"\because" D=m and D'=1

The auxiliary equation is;

m3-3m-2=0

(m+1)(m+1)(m-2)=0

m=-1,-1,2

Therefore C.F=f1(y-x)+xf2(y-x)+f3(y+2x)

Now,

P.I="\frac{1}{D^{3}-3DD'^{2}-2D'^{3}}Cos(x+2y)"

But D2=-1,D'2=-22=-4

P.I="\frac{1}{(-1)D-3D(-4)-2(-4)D'}Cos(x+2y)"

="\frac{1}{-D+12D+8D'}Cos(x+2y)"

= "\frac{1}{11D+8D'}Cos(x+2y)"

="\frac{D}{11D^{2}+2DD'}Cos(x+2y)"

="\frac{D}{11(-1)+8(-2)}Cos(x+2y)\ \because D^{2}=-1^{2}\ and\ D-D'=-1*2=-2"

P.I="\frac{D}{-11-16}\int Cos(x+2y)"

="\frac{-1}{27}[-Sin(x+2y)]"

="\frac{1}{27}Sin(x+2y)"

The general solution becomes;

y=C.F+P.I

=f1(y-x)+xf2(y-x)+f3(y+2x)+"\frac{1}{27}Sin(x+2y)"